1. At absolute zero temperature, all states within the Fermi sphere are occupied by electrons with spin state of half. Take density  cm-3. Calculate


i. Fermi wave vector.

(3)

ii. Fermi energy.

(3)

iii. Fermi momentum.



1
Expert's answer
2021-06-29T10:00:48-0400

Solution.

"n\approx10^{22}cm^{-3}=10^{28}m^{-3};"

"n=\dfrac{k_F^3}{3\pi^2}\implies k_F=(3\pi^2n)^{1/3};"

"k_F=(3\sdot3.14^2\sdot10^{28})^{1/3}=6.53\sdot10^9m^{-1};"

"E_F=\dfrac{h^2k_F^2}{2m};"

"E_F=\dfrac{(6.626\sdot10^{-34})^2\sdot(6.53\sdot10^9)^2}{2\sdot9.1\sdot10^{-31}}=15.75\sdot10^{-19}J;"

"p_F=hk_F;"

"p_F=6.626\sdot10^{-34}\sdot6.53\sdot10^9=43.27\sdot 10^{-25}kgm/s;"

Answer: I."k_F=(3\sdot3.14^2\sdot10^{28})^{1/3}=6.53\sdot10^9m^{-1};"

II."E_F=15.75\sdot10^{-19}J;"

III."p_F==43.27\sdot 10^{-25}kgm/s."